[Hallc_running] RC Daily Update

Michael Paolone tue81317 at temple.edu
Thu Feb 28 15:36:28 EST 2019

Hi Everyone,

No new major issues to report.  We have been running in our phase-2
kinematic setting since yesterday.  Beam was unstable over most of the
night, but since day shift it has been of reasonable quality at 50 to
60uA.  We plan to have a short access around 9am tomorrow to change
spectrometer settings for phase-3 running.

More info at the daily 4:30 meeting:

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190228/2e9f2327/attachment.html>

More information about the Hallc_running mailing list